OpenStudy (anonymous):

A farmer builds a rectangular pen with a perimeter of 116 feet. If the length is 5 feet more than the width, what is the width? A. w + 5 = 116 B. 4w + 10 = 116 C. 2w + 5 = 116 D. 2w + 10 = 116

OpenStudy (anonymous):

i can help

OpenStudy (anonymous):

ok yes plzz

OpenStudy (blake57roger):

set up equation w+w+5+w+w+5=116 combine 4w+10=116

OpenStudy (anonymous):

4w+10=116

OpenStudy (anonymous):

its B @courtney-ray

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

4w+10=116 4w=116-10 w=106/4 w=26.5 Now check26.5*2 +(26.5+5)*2 =53+31.5*2 =53+63 =116

OpenStudy (anonymous):

fan and medal lol:)

OpenStudy (blake57roger):

b

OpenStudy (anonymous):

HA HA i got a Medaallllllllllllllll <3

OpenStudy (anonymous):

fan and medal lets go 3

OpenStudy (anonymous):

lol u got one @openstudyhelp98765432

OpenStudy (anonymous):

yeplol

OpenStudy (anonymous):

2